# Do we still need Human Annotators? Prompting Large Language Models for Aspect Sentiment Quad Prediction

### v2

@@ E.1 TASD
Aspect Term - 0 51.80 47.39 57.23 54.53 53.23 55.98 57.86 56.47 59.72 30.09 26.14 35.46 52.79 53.45 52.22
10 56.98 53.66 61.00 61.41 58.23 65.07 49.48 56.99 43.79 43.69 39.85 48.73 52.03 57.22 47.78
20 61.04 57.29 65.39 63.21 61.08 65.59 53.83 55.18 52.56 52.66 48.03 58.35 54.10 50.46 58.40
30 63.05 59.72 66.79 66.51 65.71 67.35 56.92 55.01 59.00 53.45 48.25 60.14 56.37 54.36 58.58
40 63.97 62.27 65.79 67.00 65.76 68.30 60.03 57.01 63.46 59.10 54.13 65.22 56.64 55.27 58.12
50 63.68 61.94 65.57 67.42 64.96 70.16 57.83 52.95 63.79 56.36 51.85 61.79 58.59 57.15 60.14
SC 0 51.70 51.56 51.85 53.44 55.56 51.47 58.81 61.72 56.16 32.47 30.77 34.36 55.25 60.88 50.57
10 56.07 76.43 44.28 61.19 82.32 48.69 48.51 79.89 34.83 31.23 77.03 19.59 49.93 71.19 38.44
20 67.25 83.15 56.46 67.58 83.61 56.70 56.04 75.50 44.55 40.41 79.00 27.15 56.51 79.92 43.71
30 66.16 81.40 55.72 69.41 86.76 57.84 60.50 79.54 48.82 44.84 83.96 30.58 58.52 77.15 47.14
40 67.67 81.35 57.93 69.47 83.83 59.31 66.21 78.76 57.11 55.56 85.11 41.24 57.14 76.05 45.77
50 67.94 79.65 59.23 70.72 82.71 61.76 64.62 74.01 57.35 53.05 83.70 38.83 59.55 77.09 48.51
Aspect Category - 0 41.03 46.58 36.70 60.47 67.13 55.05 73.39 79.74 68.20 43.10 48.73 38.64 53.53 65.45 45.31
10 55.53 54.03 57.16 55.70 51.49 60.70 59.92 64.48 55.96 25.76 24.75 26.86 54.34 58.22 50.95
20 60.05 59.24 60.90 62.61 56.89 69.64 64.95 64.21 65.72 28.54 26.81 30.54 50.65 49.57 51.79
30 60.89 59.77 62.06 67.93 63.33 73.24 69.49 67.66 71.44 28.05 25.70 30.91 56.29 55.72 56.89
40 64.12 62.83 65.47 66.81 63.73 70.20 70.68 68.22 73.36 31.91 29.29 35.08 58.69 59.15 58.26
50 63.83 63.94 63.74 67.86 65.03 70.96 70.80 67.31 74.72 31.01 28.57 33.93 59.14 59.68 58.63
SC 0 37.11 45.99 31.10 59.72 69.98 52.09 72.81 84.43 64.00 42.08 52.47 35.12 52.66 70.39 42.07
10 53.69 80.05 40.39 54.24 77.50 41.72 56.56 89.22 41.40 15.46 37.90 9.71 58.53 82.87 45.24
20 60.61 82.96 47.74 64.56 78.05 55.05 63.38 85.96 50.20 20.92 44.90 13.64 55.01 82.61 41.24
30 58.90 78.96 46.97 68.70 85.09 57.60 66.00 88.55 52.60 25.04 47.95 16.94 60.17 83.88 46.91
40 63.08 79.84 52.13 66.29 81.82 55.72 70.11 87.69 58.40 28.28 46.89 20.25 59.30 81.58 46.58
50 65.45 84.49 53.42 67.31 82.87 56.66 69.95 84.66 59.60 25.32 41.71 18.18 62.05 85.84 48.58
Sentiment Polarity - 0 85.30 90.82 80.43 86.10 90.06 82.48 89.32 91.27 87.48 83.65 88.73 79.12 89.48 93.67 85.66
10 88.31 89.87 86.81 87.18 87.35 87.02 86.28 89.06 83.67 86.26 84.46 88.15 85.94 91.99 80.66
20 88.52 89.94 87.15 88.17 86.72 89.68 84.35 83.31 85.43 88.81 88.18 89.46 90.69 91.85 89.58
30 87.70 88.12 87.28 88.05 86.55 89.62 84.61 83.27 86.01 87.02 85.10 89.02 90.67 92.29 89.10
40 88.82 89.54 88.11 88.77 87.51 90.06 87.27 85.80 88.80 87.85 87.00 88.73 90.97 92.12 89.86
50 87.81 89.33 86.35 89.86 88.74 91.02 87.87 85.66 90.22 88.35 87.60 89.12 91.46 92.78 90.19
SC 0 81.69 91.88 73.53 84.54 91.50 78.57 87.92 92.68 83.62 80.00 91.25 71.22 85.60 95.10 77.83
10 68.00 96.58 52.48 67.77 96.48 52.22 67.19 95.50 51.83 41.97 93.28 27.07 76.79 97.81 63.21
20 72.76 94.55 59.13 77.29 95.13 65.08 74.85 94.76 61.86 47.46 92.25 31.95 74.45 97.70 60.14
30 74.10 96.29 60.22 78.12 95.63 66.03 75.52 96.93 61.86 53.40 93.87 37.32 78.09 96.53 65.57
40 78.31 96.38 65.94 78.39 95.44 66.51 79.20 95.83 67.48 59.63 95.19 43.41 79.05 96.92 66.75
50 76.98 95.00 64.71 79.33 96.58 67.30 83.08 97.06 72.62 61.49 95.38 45.37 79.72 97.94 67.22
@@
Table 7: Gemma-2-9B: Performance scores at element-level for the TASD task. The best score achieved by a method is presented in bold.
E.2 ASQP
@@
Aspect Term - 0 65.60 57.02 77.23 71.30 64.43 79.82 64.69 59.45 70.95 54.57 43.54 73.06 66.79 59.18 76.66
10 67.38 60.12 76.64 74.53 67.96 82.52 66.65 63.79 69.81 63.83 54.41 77.18 67.75 63.78 72.27
20 68.59 63.37 74.78 74.87 68.39 82.73 67.44 63.48 71.94 67.31 58.83 78.69 68.24 62.44 75.24
30 67.94 62.89 73.87 75.46 69.60 82.41 69.54 64.86 74.98 68.21 59.23 80.41 69.65 64.41 75.84
40 68.27 63.97 73.20 76.19 71.46 81.60 69.32 64.84 74.45 68.19 60.18 78.69 67.61 63.59 72.17
50 69.42 64.93 74.58 75.93 72.80 79.36 69.57 64.89 74.98 70.31 63.17 79.31 68.17 66.55 69.89
SC 0 67.57 62.23 73.91 74.19 71.17 77.48 66.97 64.68 69.43 59.05 51.96 68.38 68.78 63.80 74.60
10 70.68 68.39 73.12 75.94 76.34 75.53 67.33 71.05 63.98 62.25 64.93 59.79 67.81 72.89 63.39
20 69.35 70.55 68.18 77.37 76.56 78.19 69.48 72.92 66.35 66.42 72.18 61.51 69.68 74.80 65.22
30 69.24 71.19 67.39 76.62 76.83 76.42 71.27 74.29 68.48 65.91 73.42 59.79 67.72 72.09 63.84
40 70.15 72.92 67.59 77.76 79.34 76.24 71.36 74.48 68.48 65.23 76.15 57.04 66.83 74.09 60.87
50 70.53 75.45 66.21 75.66 80.85 71.10 72.07 76.05 68.48 67.98 79.36 59.45 66.14 78.37 57.21
Opinion Term - 0 62.52 60.46 64.73 65.57 64.22 66.98 57.47 58.07 56.90 29.95 26.72 34.07 44.13 46.48 42.01
10 62.29 58.89 66.10 67.53 64.50 70.87 58.60 58.13 59.08 41.73 38.03 46.24 47.95 50.52 45.64
20 56.77 55.03 58.63 68.57 66.23 71.09 58.06 57.30 58.84 43.16 39.66 47.33 50.20 51.05 49.39
30 59.05 57.00 61.26 67.86 65.79 70.08 61.24 60.46 62.05 44.18 40.78 48.19 53.09 54.07 52.15
40 59.42 56.49 62.69 67.96 66.31 69.70 59.66 58.87 60.49 43.79 40.65 47.47 51.42 52.35 50.52
50 59.74 57.19 62.53 67.93 66.48 69.46 62.21 60.84 63.63 43.10 39.80 47.01 52.02 54.29 49.94
SC 0 62.76 64.45 61.16 65.73 68.18 63.45 58.03 62.05 54.50 29.09 28.99 29.19 44.69 50.00 40.41
10 63.51 67.01 60.35 67.05 70.91 63.59 59.41 66.89 53.44 41.79 49.84 35.97 48.32 59.45 40.70
20 55.66 61.94 50.54 67.54 72.36 63.32 58.01 65.56 52.03 43.05 52.61 36.43 52.50 64.35 44.33
30 60.01 66.67 54.57 67.40 72.76 62.77 61.90 70.75 55.03 43.57 54.21 36.43 53.40 64.34 45.64
40 60.41 67.05 54.97 67.78 72.93 63.32 61.67 69.87 55.20 40.23 53.16 32.35 52.42 64.74 44.04
50 58.05 65.87 51.88 67.21 74.22 61.41 61.80 70.18 55.20 42.92 54.74 35.29 52.63 68.13 42.88
Aspect Category - 0 59.61 59.24 60.00 66.46 66.11 66.82 74.33 73.72 74.96 42.14 39.67 44.96 64.63 65.06 64.21
10 72.31 71.55 73.09 70.99 69.51 72.53 76.01 75.11 76.96 40.05 38.39 41.86 64.58 66.58 62.70
20 72.41 73.00 71.83 77.21 76.69 77.75 77.30 75.31 79.40 44.53 43.17 45.99 67.07 67.89 66.28
30 74.19 74.79 73.61 79.82 79.81 79.82 79.18 77.78 80.64 47.28 46.30 48.31 70.20 71.47 68.98
40 75.51 75.37 75.64 80.86 81.21 80.53 79.20 77.56 80.92 47.06 46.03 48.14 70.32 71.78 68.91
50 75.54 75.82 75.27 81.03 81.66 80.41 78.87 77.07 80.76 49.09 48.30 49.92 69.88 73.07 66.98
SC 0 57.86 61.55 54.58 64.35 68.03 61.05 72.63 75.70 69.80 43.33 46.05 40.91 64.96 69.11 61.27
10 71.23 78.92 64.90 69.91 75.42 65.15 72.99 81.33 66.20 41.45 52.37 34.30 63.84 75.57 55.26
20 69.93 81.37 61.32 78.05 85.79 71.60 75.00 82.14 69.00 45.04 58.61 36.57 66.41 79.35 57.10
30 73.41 85.23 64.47 79.94 88.89 72.62 76.16 84.98 69.00 47.47 63.76 37.81 67.12 80.09 57.76
40 72.98 84.82 64.04 80.64 89.32 73.50 76.81 84.78 70.20 42.44 59.26 33.06 67.85 82.89 57.43
50 73.53 88.51 62.89 79.67 90.04 71.45 76.63 85.86 69.20 46.56 64.71 36.36 66.94 84.87 55.26
Sentiment Polarity - 0 90.19 90.02 90.36 90.99 90.54 91.45 89.52 90.30 88.75 88.09 87.60 88.59 90.13 90.85 89.43
10 90.41 89.89 90.94 91.68 90.86 92.52 90.45 91.49 89.44 87.04 87.07 87.02 86.88 91.19 82.97
20 90.34 89.74 90.94 92.58 92.71 92.45 91.79 91.46 92.13 88.84 89.00 88.68 89.50 91.40 87.69
30 90.79 90.78 90.80 93.23 93.42 93.03 93.02 92.51 93.55 88.99 89.07 88.93 90.33 93.14 87.69
40 91.56 91.16 91.97 94.10 94.17 94.03 92.88 92.03 93.74 89.46 89.11 89.80 89.97 92.27 87.78
50 91.38 91.24 91.52 93.41 93.69 93.14 93.28 92.54 94.03 89.34 89.36 89.32 90.84 94.36 87.59
SC 0 86.92 91.34 82.91 88.75 92.84 85.00 87.26 92.12 82.89 81.15 87.57 75.61 87.95 92.91 83.49
10 85.66 92.64 79.66 86.62 93.95 80.34 84.64 94.29 76.77 72.14 90.44 60.00 82.38 93.96 73.35
20 83.14 93.21 75.04 87.02 94.91 80.34 85.03 93.81 77.75 74.45 94.38 61.46 81.50 94.41 71.70
30 84.04 93.88 76.07 88.01 96.31 81.03 84.71 94.85 76.53 70.87 92.19 57.56 81.63 96.46 70.75
40 83.62 94.41 75.04 87.62 96.09 80.52 86.13 94.72 78.97 70.73 94.31 56.59 81.78 95.58 71.46
50 82.01 94.43 72.48 85.96 95.57 78.10 85.79 96.06 77.51 70.52 93.55 56.59 80.72 97.02 69.10
@@
Table 8: Gemma-2-27B: Performance scores at element-level for the ASQP task. The best score achieved by a method is presented in bold.

### v3

@@ E.1 ASQP
Aspect Term - 0 42.65 40.10 45.57 47.24 46.76 47.73 41.88 44.14 39.86 42.38 38.79 46.74 48.45 52.61 44.94
10 44.72 51.77 39.37 43.20 60.32 33.69 46.49 57.73 38.96 58.02 60.39 55.88 50.93 66.01 41.56
20 54.07 63.55 47.08 56.39 68.74 47.84 46.03 59.07 37.73 61.45 66.57 57.11 56.04 70.06 46.73
30 53.00 63.18 45.69 57.79 72.11 48.26 46.91 59.23 38.86 63.85 69.52 59.11 54.14 71.10 43.75
40 50.04 64.61 40.87 55.96 68.92 47.13 48.45 61.49 40.00 66.97 71.39 63.09 59.82 75.44 49.61
50 56.24 67.42 48.26 55.65 71.51 45.60 51.23 61.49 43.93 68.95 74.19 64.40 59.81 76.22 49.24
SC 0 21.48 71.11 12.65 20.03 69.47 11.70 28.19 76.04 17.30 18.56 72.09 10.65 27.60 79.35 16.70
10 12.48 87.18 6.72 13.07 83.33 7.09 15.42 80.00 8.53 12.74 86.96 6.87 19.55 88.89 10.98
20 31.07 85.71 18.97 34.51 85.31 21.63 14.29 82.50 7.82 23.88 90.91 13.75 31.95 89.47 19.45
30 33.85 78.99 21.54 40.71 86.71 26.60 23.69 77.63 13.98 29.15 96.15 17.18 26.80 88.46 15.79
40 32.75 83.74 20.36 46.57 86.12 31.91 27.95 82.56 16.82 36.41 87.01 23.02 39.86 92.50 25.40
50 44.57 83.24 30.43 42.91 84.82 28.72 40.93 82.14 27.25 38.93 86.90 25.09 40.58 97.39 25.63
Opinion Term - 0 30.92 33.74 28.55 31.50 34.56 28.94 25.50 30.46 21.94 19.31 20.36 18.37 23.71 31.51 19.01
10 34.53 39.54 30.65 37.36 41.95 33.70 30.56 37.51 25.82 19.77 19.99 19.55 22.96 28.93 19.04
20 38.89 41.66 36.48 46.82 50.41 43.72 31.00 36.42 26.98 24.05 23.29 24.89 30.04 33.62 27.15
30 43.28 45.24 41.51 48.23 52.42 44.67 36.58 42.66 32.03 27.39 26.89 27.92 27.15 36.02 21.80
40 44.38 45.72 43.12 52.42 54.53 50.49 36.23 42.56 31.53 27.65 29.06 26.38 29.80 39.22 24.04
50 48.43 50.02 46.94 52.56 55.77 49.70 43.89 49.32 39.54 26.81 28.96 24.98 32.82 42.88 26.60
SC 0 15.73 69.47 8.87 15.38 66.67 8.70 17.77 60.82 10.41 7.44 42.86 4.07 14.81 61.05 8.43
10 13.53 79.71 7.39 17.61 78.49 9.92 14.17 81.48 7.76 12.11 78.38 6.56 12.58 64.00 6.98
20 27.59 67.54 17.34 34.30 81.22 21.74 14.06 74.58 7.76 15.63 68.42 8.82 20.34 60.87 12.21
30 34.91 73.91 22.85 38.27 78.81 25.27 22.97 69.64 13.76 17.73 59.74 10.41 15.74 51.20 9.30
40 36.94 76.69 24.33 46.27 75.85 33.29 27.51 73.28 16.93 19.81 62.65 11.76 19.64 59.12 11.77
50 43.79 74.28 31.05 45.01 74.76 32.20 39.47 77.72 26.46 22.01 62.77 13.35 21.93 58.13 13.52
Aspect Category - 0 37.64 42.23 33.95 39.84 44.53 36.05 51.54 62.70 43.76 29.56 34.72 25.74 42.32 54.98 34.42
10 45.75 52.79 40.37 52.52 57.49 48.35 55.00 64.50 47.96 37.10 41.08 33.84 54.40 67.51 45.58
20 58.32 62.95 54.33 60.45 64.79 56.66 59.89 67.80 53.64 46.24 49.09 43.72 58.27 67.28 51.39
30 62.24 66.00 58.88 66.41 71.72 61.84 63.17 71.11 56.84 43.71 45.83 41.78 61.57 68.68 55.79
40 65.14 68.71 61.92 68.52 73.06 64.51 67.03 74.69 60.80 42.98 45.90 40.41 63.61 69.91 58.36
50 67.82 71.86 64.21 68.79 74.81 63.66 68.63 74.52 63.60 42.88 46.89 39.50 61.46 70.48 54.49
SC 0 13.13 55.32 7.45 15.42 63.16 8.78 26.89 84.21 16.00 10.23 61.36 5.58 23.38 86.17 13.52
10 13.11 76.92 7.16 20.13 84.78 11.42 17.30 87.27 9.60 11.13 78.38 5.99 19.94 91.78 11.19
20 35.27 85.64 22.21 36.11 82.29 23.13 18.60 88.14 10.40 16.70 81.82 9.30 31.34 85.19 19.20
30 40.83 87.32 26.65 43.83 88.44 29.14 31.48 87.27 19.20 18.67 71.23 10.74 37.43 86.67 23.87
40 41.82 85.78 27.65 52.70 86.33 37.92 36.51 88.46 23.00 22.22 69.57 13.22 43.80 90.58 28.88
50 50.25 86.67 35.39 51.02 84.18 36.60 47.35 89.44 32.20 24.37 67.29 14.88 39.33 85.47 25.54
Sentiment Polarity - 0 77.70 83.22 72.89 76.66 82.89 71.31 74.38 83.45 67.09 71.14 81.44 63.17 74.19 83.92 66.51
10 77.94 85.50 71.62 80.40 84.25 76.90 77.53 83.91 72.08 78.14 82.50 74.24 78.03 86.94 70.80
20 86.28 89.08 83.66 87.12 89.61 84.76 80.61 85.23 76.48 84.19 85.93 82.54 83.29 90.03 77.50
30 87.57 90.86 84.51 88.53 92.13 85.21 81.46 85.66 77.65 82.60 84.60 80.68 87.91 91.92 84.25
40 88.28 92.00 84.85 89.57 93.21 86.21 83.02 88.82 77.95 84.77 88.48 81.37 90.27 94.20 86.65
50 90.00 93.32 86.91 90.06 94.24 86.24 85.28 89.01 81.86 86.57 90.97 82.59 88.85 94.23 84.06
SC 0 26.33 97.80 15.21 25.85 93.55 15.00 37.30 98.95 22.98 17.70 95.24 9.76 34.88 97.83 21.23
10 18.77 93.85 10.43 25.07 93.33 14.48 22.89 98.15 12.96 15.66 94.59 8.54 27.59 98.55 16.04
20 44.88 96.61 29.23 46.27 95.68 30.52 25.21 100.00 14.43 21.55 92.59 12.20 45.05 95.42 29.48
30 50.38 97.07 34.02 53.02 97.69 36.38 41.31 98.17 26.16 27.80 93.06 16.34 51.55 94.94 35.38
40 51.32 97.14 34.87 64.19 97.88 47.76 46.64 98.43 30.56 33.60 93.33 20.49 57.10 97.71 40.33
50 62.00 97.44 45.47 63.49 97.50 47.07 57.29 95.98 40.83 38.13 94.23 23.90 54.15 95.81 37.74
@@
Table 7: Gemma-3-4B: Performance scores at element-level for the ASQP task. The best score achieved with respect to a sentiment element is presented in bold.
E.2 TASD
@@
Aspect Term - 0 60.81 49.78 78.12 68.31 58.42 82.22 59.63 50.21 73.41 53.48 40.24 79.73 61.69 51.08 77.85
10 73.78 67.50 81.37 78.07 74.59 81.90 70.06 65.93 74.74 67.85 58.58 80.62 74.36 69.41 80.09
20 74.38 68.13 81.88 79.88 76.42 83.66 69.52 66.17 73.22 72.83 66.57 80.41 75.68 70.35 81.88
30 75.50 74.29 76.75 77.98 75.74 80.36 71.41 69.07 73.93 73.84 70.11 78.01 75.42 73.19 77.80
40 74.58 75.46 73.73 77.36 76.86 77.88 70.51 69.73 71.33 72.74 72.58 72.92 74.98 74.83 75.15
50 74.09 73.90 74.28 76.40 75.02 77.84 70.55 66.71 74.88 70.82 70.72 71.00 74.22 74.50 73.96
SC 0 61.28 50.30 78.41 68.49 58.78 82.03 59.85 50.49 73.46 53.89 40.70 79.73 61.61 51.29 77.12
10 74.47 68.92 81.00 78.36 75.57 81.37 70.71 67.16 74.64 68.87 61.23 78.69 75.00 70.88 79.63
20 75.63 69.84 82.47 80.13 77.16 83.33 70.18 68.00 72.51 73.23 69.00 78.01 76.32 72.34 80.78
30 76.13 76.06 76.20 78.16 77.11 79.25 73.24 72.56 73.93 76.82 77.35 76.29 75.35 75.87 74.83
40 75.23 76.53 73.99 78.42 79.07 77.78 73.27 73.80 72.75 73.70 76.87 70.79 77.01 78.57 75.51
50 76.25 78.25 74.35 77.97 78.16 77.78 73.85 71.56 76.30 72.93 78.57 68.04 76.61 80.05 73.46
Aspect Category - 0 56.36 56.46 56.26 72.40 71.76 73.06 82.79 79.14 86.80 49.68 47.86 51.65 67.23 66.23 68.25
10 72.09 74.06 70.22 81.87 81.68 82.07 82.49 81.36 83.64 50.39 51.04 49.75 70.42 71.91 68.98
20 75.99 79.55 72.75 81.52 82.57 80.48 82.08 81.61 82.56 56.17 57.95 54.50 70.32 71.85 68.85
30 78.30 82.86 74.22 83.17 84.46 81.91 81.66 82.29 81.04 56.58 58.16 55.08 74.76 77.29 72.39
40 79.63 84.52 75.28 83.74 85.62 81.94 81.61 82.86 80.40 54.72 57.21 52.44 75.01 78.98 71.42
50 79.19 83.89 74.99 83.74 85.69 81.88 81.03 81.34 80.72 53.54 56.50 50.87 74.50 79.22 70.32
SC 0 56.05 56.23 55.87 72.19 71.71 72.68 82.90 79.34 86.80 49.95 48.36 51.65 67.49 66.89 68.11
10 72.98 76.16 70.06 82.44 83.06 81.83 82.59 82.18 83.00 51.08 53.64 48.76 71.55 73.98 69.28
20 75.72 80.32 71.61 81.57 83.40 79.81 81.85 82.52 81.20 58.54 63.16 54.55 70.61 73.68 67.78
30 78.39 84.60 73.03 83.60 86.81 80.62 80.70 83.37 78.20 58.57 64.99 53.31 75.20 80.42 70.62
40 80.08 86.29 74.71 83.79 87.54 80.35 82.04 85.31 79.00 56.13 62.98 50.62 75.81 81.82 70.62
50 79.02 86.62 72.65 84.51 88.63 80.75 80.99 83.30 78.80 55.14 63.44 48.76 76.12 84.21 69.45
Sentiment Polarity - 0 87.13 88.58 85.73 89.63 90.49 88.79 92.21 90.18 94.33 87.22 86.55 87.90 90.32 90.21 90.42
10 91.81 93.29 90.37 93.00 94.35 91.68 94.77 94.33 95.21 90.69 91.06 90.34 92.33 93.26 91.42
20 92.01 93.91 90.19 93.73 95.01 92.48 93.92 93.76 94.08 91.16 91.79 90.54 93.34 94.45 92.26
30 92.60 94.87 90.43 93.79 95.02 92.60 93.46 93.58 93.35 90.40 91.46 89.37 93.20 94.65 91.79
40 92.63 95.14 90.25 93.19 94.81 91.62 94.04 94.35 93.74 90.38 92.10 88.73 93.12 95.00 91.32
50 92.58 95.10 90.19 93.74 94.86 92.63 93.17 93.33 93.01 89.62 91.78 87.56 93.25 95.36 91.23
SC 0 87.31 88.92 85.76 89.89 90.91 88.89 92.09 90.35 93.89 87.48 87.17 87.80 90.33 90.33 90.33
10 91.43 93.81 89.16 93.13 94.89 91.43 94.23 94.58 93.89 89.53 92.69 86.59 91.98 93.43 90.57
20 91.90 94.30 89.63 93.28 95.21 91.43 93.33 94.26 92.42 88.27 92.51 84.39 92.62 95.04 90.33
30 91.23 94.97 87.77 93.47 96.14 90.95 92.27 94.15 90.46 86.61 92.76 81.22 92.48 95.25 89.86
40 91.87 95.64 88.39 92.04 95.25 89.05 92.60 95.10 90.22 86.13 92.94 80.24 93.22 95.77 90.80
50 91.34 95.76 87.31 93.05 95.95 90.32 93.38 95.41 91.44 85.03 93.04 78.29 90.68 95.80 86.08
@@
Table 8: Gemma-3-27B: Performance scores at element-level for the TASD task. The best score achieved by a method is presented in bold.